# Convergence and Optimality of the EM Algorithm Under Multi-Component Gaussian Mixture Models

### v1

@@ 4.4 Application to community detection
  • If $\Delta_{\max}\leq\log n$, assume (4.8) with $\widehat{\theta}^{(0)}$ satisfying either (3.5) or (4.16);
  • If $\Delta_{\max}\geq\log n$, assume (4.14) with $\widehat{\theta}^{(0)}$ satisfying either (4.15) or (4.16);
@@
It is worth comparing Theorem 16 with the existing literature on the community detection problem under GMMs. A large body of work focuses on exact label recovery, with spectral clustering being the primary driving approach [12, 37, 41, 27, 3, 1]. Along this direction, the best known separation condition for exact recovery [1] requires
@@
provided that $n\pi_{\min}^{*}\gtrsim L(d+\log L)\log n$. By contrast, Theorem 16 for exact recovery requires the separation condition $\Delta_{\min}\gtrsim\log(1/\pi_{\min}^{*})+\log n$. Since $\Delta_{\min}\geq\min_{k\neq\ell}\|\mu_{k}^{*}-\mu_{\ell}^{*}\|_{2}^{2}/\lambda_{1}(\Sigma^{*})$, ours is weaker than (4.22), especially when $\pi_{\min}^{*}$ is small or $L$ large. As shown in [35], the requirement $\Delta_{\min}\gtrsim\log n$ is necessary for exact recovery even for the symmetric, isotropic 2-GMM. Regarding the sample complexity, under condition (4.22), our condition (4.14) simplifies to $dL^{2}\log(n\Delta_{\max})\lesssim n$ which is milder when $\log(\Delta_{\max}/\Delta_{\min})\ll 1/\pi_{\min}^{*}$.
The first rate of convergence for the misclustering error was established in Lu and Zhou [30], where the authors showed that Lloyd’s algorithm, when initialized by spectral clustering and run for $\lceil 4\log n\rceil$ iterations, yields estimated labels $\widehat{Y}_{1},\ldots,\widehat{Y}_{n}$ that satisfy
@@
under the assumption that $L\log n\leq n\pi_{\min}^{*2}$. In comparison, Theorem 16 requires a weaker separation when $\min\{\log(\Delta_{\max}/\Delta_{\min}),\log\log n\}\ll L/\pi_{\min}^{*}$. This is due to the fact that the EM algorithm uses soft assignments in the E-steps, which influence the updates of the mean vectors in the M-steps. On the other hand, under condition (4.24), Theorem 16 still requires $dL\log^{4}(n)\leq n\pi_{\min}^{*}$, which is a stronger requirement on the sample size $n$ than that in [30]. This is attributable to the fact that the EM algorithm also estimates the covariance matrix $\Sigma^{*}$.
Precisely because Lloyd’s algorithm does not explicitly estimate $\Sigma^{*}$, the rates in (4.23) are only optimal in the special case where $\Sigma^{*}\propto\mathbf{I}_{d}$. Chen and Zhang [11] recently established the minimax lower bound of the misclustering error for a fixed, general $\Sigma^{*}$ under the $L$-GMM:
@@
provided that $\Delta_{\min}/\log L\to\infty$. Here $\mathbb{E}_{X\mid Y}$ denotes the conditional expectation of $X_{1},\ldots,X_{n}$ under model (1.1) when $Y_{1},\ldots,Y_{n}$ are treated as deterministic quantities. In [11], they also show that a modified Lloyd’s algorithm that uses the hard assignments but also estimates $\Sigma^{*}$ provably achieves the above lower bound under conditions including $L=\mathcal{O}(1)$, $1/\pi_{\min}^{*}=\mathcal{O}(1)$, $d^{2}=\mathcal{O}(n)$, $\Delta_{\min}\to\infty$ and $c\leq\lambda_{d}(\Sigma^{*})\leq\lambda_{1}(\Sigma^{*})\leq C$, as well as an initialization such that $\sum_{i=1}^{n}1\{Y_{i}\neq\widehat{Y}^{(0)}\}=o(n)$. By contrast, our result in Theorem 16 shows that the EM algorithm also achieves the optimal rate of convergence under milder conditions, owing to the use of soft assignments.
@@ 5.1 Choice of the initialization
Our analysis in Sections 3 and 4 has focused on the convergent behavior of EM under the assumption that a suitable initialization is provided. In this section, we discuss practical procedures for choosing such initializations. Recall that our theoretical results accommodate two types of initialization schemes: (a) Initializing with $\theta^{(0)}$ that is close to $\theta^{*}$ in the sense of (3.5), or (b) Initializing the posterior probabilities $\gamma_{\ell}(X_{i};\theta^{(0)})$ for $i\in[n]$ and $\ell\in[L]$ such that $\phi(\theta^{(0)})$ in (4.10) is small, as required by (4.16).
@@
This leads to sufficient conditions on the proportion of misclassified data points under which (4.13) and (4.16) hold. For example, when the separation condition (4.24) is satisfied, one can use the Lloyd’s algorithm to obtain preliminary labels. Indeed, the rates of Lloyd’s algorithm given in (4.23) ensure that both (4.13) and (4.16) hold, provided the following balancing condition is met:
@@
On the other hand, when the separation between components is insufficient, scheme (a) can still be employed, provided that a consistent estimator of $\theta^{*}$ is available. In the context of Remark 4, it can be further simplified to find a consistent estimator of $\pi^{*}$ and $M^{*}$. A consistent estimator of $\theta^{*}$ can be obtained, for example, via the Method of Moments (MoM) approaches mentioned in the Introduction. When $d=1$, the denoising MoM estimator proposed in [45] achieves a convergence rate of $(1/\pi_{\min}^{*}+1/\Delta_{\min})n^{-1/\mathcal{O}(L)}$. This exponential dependence on $L$ is unavoidable in the absence of any separation condition, and can be improved as the separation increases. For the multivariate $L$-GMM, a fast and consistent MoM estimation is proposed in [29], though no explicit convergence rates are provided. Recently, for higher dimensions ($d\geq 2$) and $L=2$, Kalai et al. [26] proposed a MoM procedure that provably recovers the mixture parameters under the minimal separation. This framework was later extended to $L\geq 3$ by Moitra and Valiant [34]. While their estimation errors exhibit only polynomial dependence on $d$, $1/\pi_{\min}^{*}$, and $1/\Delta_{\min}$, the degree of such polynomial is high, and the required sample complexity still grows exponentially in $L$. More recently, Doss et al. [16] developed an estimator for the mixing measure that is minimax optimal in the multivariate isotropic $L$-GMM setting. In principle, this estimator could be used to construct a consistent estimator of $\theta^{*}$ with lower-order dependence on $d$, but this direction is not pursued in their work. For general anisotropic $L$-GMMs with $L\geq 3$, designing a provably consistent and computationally efficient MoM estimator with low sample complexity remains an open and important direction for future research.
@@ Appendix A A full comparison table with existing theoretical works of the EM algorithm under the $L$-GMM
Table 3: Comparison with existing theoretical results for the EM algorithm under the $L$-GMM. Symmetric 2-GMM means $\pi_{1}=\pi_{2}=1/2$ and $\mu_{1}=-\mu_{2}=\mu$. In [9], $\beta=\Sigma^{-1}(\mu_{2}-\mu_{1})$.
@@
[4] 22 symmetric, $\Sigma=\mathbf{I}_{d}$ >C>C $\|\mu^{(0)}-\mu\|_{2}\leq\|\mu\|_{2}/4$ $\mathcal{O}(\|\mu\|_{2}^{3}\sqrt{d/n})$
[47] 22 symmetric, $\Sigma$ known >0>0 $\mu^{(0)\top}\mu\neq 0$ $o(1)$
[14] 22 symmetric, $\Sigma$ known $\widetilde{\Omega}((d/n)^{1/3})$ $\mu^{(0)\top}\mu\neq 0$ $\widetilde{\mathcal{O}}(\sqrt{d/n}\wedge\|\mu\|_{2}^{-2}\sqrt{d/n})$
[46] 22 symmetric, $\Sigma=\mathbf{I}_{d}$ $(0,C]$ $\mu^{(0)\top}\mu\neq 0$ $\widetilde{\mathcal{O}}\left((d/n)^{1/4}\wedge\|\mu\|_{2}^{-1}\sqrt{d/n}\right)$
[43] 22
@@
$(0,C]$ $\mu^{(0)\top}\mu\geq 0$ $\widetilde{\mathcal{O}}\left((d/n)^{1/4}\wedge\|\mu\|_{2}^{-1}\sqrt{d/n}\wedge(\pi_{2}-\pi_{1})^{-1}\sqrt{d/n}\right)$
[9] 22
@@
$d(\widehat{M},M)+\|\widehat{\beta}-\beta\|_{2}=\widetilde{\mathcal{O}}(\sqrt{d/n})$
[13] $\geq 2$ $\Sigma=\sigma^{2}\mathbf{I}_{d}$ $\Omega\Bigl{(}\sqrt{d\log\pi_{\min}^{-1}}+\log\pi_{\min}^{-1}\Bigr{)}$ A particular initialization $d(\widehat{\pi},\pi)+d(\widehat{M},M)=\mathcal{O}\left(\sqrt{d\over n\pi_{\min}}+{\exp(-c\sqrt{d})\over\pi_{\min}}\right)$
[50] $\geq 2$ $\pi$ known, $\Sigma=\mathbf{I}_{d}$ $\Omega\Bigl{(}(d\wedge L)\log{\Delta_{\max}\over\pi_{\min}}\Bigr{)}$ $d(M^{(0)},M)\leq c\sqrt{\Delta_{\min}}$ $d(\widehat{M},M)=\widetilde{\mathcal{O}}\left({\pi_{\max}\over\pi_{\min}}\sqrt{\Delta_{\max}}(\sqrt{d}+L^{3}\Delta_{\max})\sqrt{d\over n}\right)$
[51] $\geq 2$ $\pi$ known, $\Sigma=\mathbf{I}_{d}$ $\Omega\Bigl{(}(d\wedge L)\log{\Delta_{\max}\over\pi_{\min}}\Bigr{)}$ $d(M^{(0)},M)\leq c\sqrt{\Delta_{\min}}$ $d(\widehat{M},M)=\widetilde{\mathcal{O}}\left({\sqrt{\Delta_{\max}}\over\pi_{\min}}\sqrt{dL\over n}\right)$
This work $\geq 2$ Full generality $\Omega\Bigl{(}\log{\Delta_{\max}\over\pi_{\min}}\Bigr{)}$ $d(M^{(0)},M)\leq c\sqrt{\Delta_{\min}}$ $d(\widehat{\pi},\pi)+d(\widehat{M},M)=\widetilde{\mathcal{O}}\Bigl{(}\sqrt{d\over n\pi_{\min}}\Bigr{)}$
$d(\pi^{(0)},\pi)<1,~d(\Sigma^{(0)},\Sigma)\leq c$,
or via preliminary labels
$d(\widehat{\Sigma},\Sigma)=\widetilde{\mathcal{O}}\left(\sqrt{d/n}\right)$
@@ C.3 Proof of Theorem 2: contraction of the population level EM iterates
$\begin{split}\mathcal{R}_{I}^{(0)}&:=\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\max_{\ell\in[L]}\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)\right]\sqrt{\Delta_{a\ell}},\\ \mathcal{R}_{I}^{(1)}&:=\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\max_{\ell\in[L]}~\Bigl{\|}\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)NN^{\top}\right]\Bigr{\|}_{\rm op},\\ \mathcal{R}_{I}^{(2)}&:=\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\max_{\ell\in[L]}\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\left\|\mathbb{E}\left[I_{a\ell}(X,\omega)NN^{\top}\right]\right\|_{\rm op},\end{split}$ (C.22)
@@
[…]
@@
For any $\omega=(\pi,M,J)$, recall that
$\bm{\mu}_{\ell}(\omega)=\frac{\mathbb{E}[\gamma_{\ell}(X;\omega)X]}{\mathbb{E}[\gamma_{\ell}(X;\omega)]},\qquad\forall~\ell\in[L].$
Basic calculation of partial derivatives of $\bm{\mu}_{\ell}(\omega)$ gives that
$\partial\bm{\mu}_{\ell}(\omega)=\frac{1}{\bm{\pi}_{\ell}(\omega)}\mathbb{E}\left[\left(X-\bm{\mu}_{\ell}(\omega)\right)\partial\gamma_{\ell}(X;\omega)^{\top}\right].$
It then follows that
@@
$\begin{split}\nabla^{\pi}\bm{\mu}_{\ell}(\bar{\omega})&:=\int_{0}^{1}{\partial\bm{\mu}_{\ell}(\omega)\over\partial\pi}\Big{|}_{\omega=(\pi_{u},\bar{M},\bar{J})}{\rm d}u\\ &=\sum_{k=1}^{L-1}\int_{0}^{1}\frac{1}{\bm{\pi}_{\ell}(\omega)}\mathbb{E}\left[\left(X-\bm{\mu}_{\ell}(\omega)\right)\left\langle{\partial\gamma_{\ell}(X;\omega)\over\partial\pi_{k}}\Big{|}_{\omega=(\pi_{u},\bar{M},\bar{J})},~\left(\bar{\pi}_{k}-\pi_{k}^{*}\right)\right\rangle\right]{\rm d}u\\ \nabla^{M}\bm{\mu}_{\ell}(\bar{\omega})&:=\int_{0}^{1}{\partial\bm{\mu}_{\ell}(\omega)\over\partial\pi}\Big{|}_{\omega=(\pi^{*},M_{u},\bar{J})}{\rm d}u\\ &=\sum_{k=1}^{L}\int_{0}^{1}\frac{1}{\bm{\pi}_{\ell}(\omega)}\mathbb{E}\left[\left(X-\bm{\mu}_{\ell}(\omega)\right)\left\langle{\partial\gamma_{\ell}(X;\omega)\over\partial\mu_{k}}\Big{|}_{\omega=(\pi^{*},M_{u},\bar{J})},~\bar{\mu}_{k}-\mu_{k}^{*}\right\rangle\right]{\rm d}u\\ \nabla^{J}\bm{\mu}_{\ell}(\bar{\omega})&:=\int_{0}^{1}{\partial\bm{\mu}_{\ell}(\omega)\over\partial\pi}\Big{|}_{\omega=(\pi^{*},M^{*},J_{u})}{\rm d}u\\ &=\int_{0}^{1}\frac{1}{\bm{\pi}_{\ell}(\omega)}\mathbb{E}\left[\left(X-\bm{\mu}_{\ell}(\omega)\right)\left\langle{\partial\gamma_{\ell}(X;\omega)\over\partial J}\Big{|}_{\omega=(\pi^{*},M^{*},J_{u})},~\bar{J}-J^{*}\right\rangle\right]{\rm d}u.\end{split}$ (C.42)
Fix any $\bar{\omega}\in\mathcal{B}_{d}(\omega^{*})$ and any $\ell\in[L]$. The main task below is to bound each terms in terms of $\mathcal{R}_{I}^{(j)}$ with $j\in\{0,1,2\}$ as well as the quantity
$D_{\mu_{\ell}}:=\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\|\bm{\mu}_{\ell}(\omega)-\bm{\mu}_{\ell}(\omega^{*})\|_{\Sigma^{*}}.$ (C.43)
For future reference, we notice that
@@
$\left\|\nabla^{\pi}\bm{\mu}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}\leq\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\frac{1}{\bm{\pi}_{\ell}(\omega)}\left\|\sum_{k=1}^{L-1}(\bar{\pi}_{k}-\pi_{k}^{*})~\mathbb{E}\left[\left(X-\bm{\mu}_{\ell}(\omega)\right){\partial\gamma_{\ell}(X;\omega)\over\partial\pi_{k}}\right]\right\|_{\Sigma^{*}}.$

### v2

@@ 4.4 Application to community detection
  • If $\Delta_{\max}\leq\log n$, assume (4.7) with $\widehat{\theta}^{(0)}$ satisfying either (3.5) or (b) of Theorem 10;
  • If $\Delta_{\max}\geq\log n$, assume (4.11) with $\widehat{\theta}^{(0)}$ satisfying either (a) or (b) of Theorem 10;
@@
It is worth comparing Theorem 15 with the existing literature on the community detection problem under GMMs. A large body of work focuses on exact label recovery, with spectral clustering being the primary driving approach [dasgupta1999learning, sanjeev2001learning, vempala2002spectral, kannan2008spectral, awasthi2012improved, achlioptas2005spectral]. Along this direction, the best known separation condition for exact recovery [achlioptas2005spectral] requires
@@
provided that $n\pi_{\min}^{*}\gtrsim L(d+\log L)\log n$. By contrast, Theorem 15 for exact recovery requires the separation condition $\Delta_{\min}\gtrsim\log(1/\pi_{\min}^{*})+\log n$. Since $\Delta_{\min}\geq\min_{k\neq\ell}\|\mu_{k}^{*}-\mu_{\ell}^{*}\|_{2}^{2}/\lambda_{1}(\Sigma^{*})$, ours is weaker than (4.17), especially when $\pi_{\min}^{*}$ is small or $L$ large. As shown in [Ndaoud2022], the requirement $\Delta_{\min}\gtrsim\log n$ is necessary for exact recovery even for the symmetric, isotropic 2-GMM. Regarding the sample complexity, under condition (4.17), our condition (4.11) simplifies to $dL^{2}\log(n\Delta_{\max})\lesssim n$ which is milder when $\log(\Delta_{\max}/\Delta_{\min})\ll 1/\pi_{\min}^{*}$.
The first rate of convergence for the misclustering error was established in lu2016statistical, where the authors showed that Lloyd’s algorithm, when initialized by spectral clustering and run for $\lceil 4\log n\rceil$ iterations, yields estimated labels $\widehat{Y}_{1},\ldots,\widehat{Y}_{n}$ that satisfy
@@
under the assumption that $L\log n\leq n\pi_{\min}^{*2}$. In comparison, Theorem 15 requires a weaker separation when $\min\{\log(\Delta_{\max}/\Delta_{\min}),\log\log n\}\ll L/\pi_{\min}^{*}$. This is due to the fact that the EM algorithm uses soft assignments in the E-steps, which influence the updates of the mean vectors in the M-steps. On the other hand, under condition (4.19), Theorem 15 still requires $dL\log^{4}(n)\leq n\pi_{\min}^{*}$, which is a stronger requirement on the sample size $n$ than that in [lu2016statistical]. This is attributable to the fact that the EM algorithm also estimates the covariance matrix $\Sigma^{*}$.
Precisely because Lloyd’s algorithm does not explicitly estimate $\Sigma^{*}$, the rates in (4.18) are only optimal in the special case where $\Sigma^{*}\propto\mathbf{I}_{d}$. chen2024achieving recently established the minimax lower bound of the misclustering error for a fixed, general $\Sigma^{*}$ under the $L$-GMM:
@@
provided that $\Delta_{\min}/\log L\to\infty$. Here $\mathbb{E}_{X\mid Y}$ denotes the conditional expectation of $X_{1},\ldots,X_{n}$ under model (1.1) when $Y_{1},\ldots,Y_{n}$ are treated as deterministic quantities. In [chen2024achieving], they also show that a modified Lloyd’s algorithm that uses the hard assignments but also estimates $\Sigma^{*}$ provably achieves the above lower bound under conditions including $L=\mathcal{O}(1)$, $1/\pi_{\min}^{*}=\mathcal{O}(1)$, $d^{2}=\mathcal{O}(n)$, $\Delta_{\min}\to\infty$ and $c\leq\lambda_{d}(\Sigma^{*})\leq\lambda_{1}(\Sigma^{*})\leq C$, as well as an initialization such that $\sum_{i=1}^{n}1\{Y_{i}\neq\widehat{Y}^{(0)}\}=o(n)$. By contrast, our result in Theorem 15 shows that the EM algorithm also achieves the optimal rate of convergence under milder conditions, owing to the use of soft assignments.
@@ 5.1 Choice of the initialization
Our analysis in Sections 3 and 4 has focused on the convergent behavior of EM under the assumption that a suitable initialization is provided. In this section, we discuss practical procedures for choosing such initializations. Recall that our theoretical results accommodate two types of initialization schemes: (a) Initializing with $\theta^{(0)}$ that is close to $\theta^{*}$ in the sense of (3.5), or (b) Initializing the posterior probabilities $\gamma_{\ell}(X_{i};\theta^{(0)})$ for $i\in[n]$ and $\ell\in[L]$ such that $\phi(\theta^{(0)})$ in (4.9) is small, as required by (b) of Theorem 10.
@@
This leads to sufficient conditions on the proportion of misclassified data points under which (b) of Theorem 9 and (b) of Theorem 10 hold. For example, when the separation condition (4.19) is satisfied, one can use the Lloyd’s algorithm to obtain preliminary labels. Indeed, the rates of Lloyd’s algorithm given in (4.18) ensure that both (b) of Theorem 9 and (b) of Theorem 10 hold, provided the following balancing condition is met:
@@
On the other hand, when the separation between components is insufficient, scheme (a) can still be employed, provided that a consistent estimator of $\theta^{*}$ is available. In the context of Remark 4, it can be further simplified to find a consistent estimator of $\pi^{*}$ and $M^{*}$. A consistent estimator of $\theta^{*}$ can be obtained, for example, via the Method of Moments (MoM) approaches mentioned in the Introduction. When $d=1$, the denoising MoM estimator proposed in [WuYang2020] achieves a convergence rate of $(1/\pi_{\min}^{*}+1/\Delta_{\min})n^{-1/\mathcal{O}(L)}$. This exponential dependence on $L$ is unavoidable in the absence of any separation condition, and can be improved as the separation increases. For the multivariate $L$-GMM, a fast and consistent MoM estimation is proposed in [LindsayBasak], though no explicit convergence rates are provided. Recently, for higher dimensions ($d\geq 2$) and $L=2$, kalai2010efficiently proposed a MoM procedure that provably recovers the mixture parameters under the minimal separation. This framework was later extended to $L\geq 3$ by moitra2010settling. While their estimation errors exhibit only polynomial dependence on $d$, $1/\pi_{\min}^{*}$, and $1/\Delta_{\min}$, the degree of such polynomial is high, and the required sample complexity still grows exponentially in $L$. More recently, doss2023optimal developed an estimator for the mixing measure that is minimax optimal in the multivariate isotropic $L$-GMM setting. In principle, this estimator could be used to construct a consistent estimator of $\theta^{*}$ with lower-order dependence on $d$, but this direction is not pursued in their work. For general anisotropic $L$-GMMs with $L\geq 3$, designing a provably consistent and computationally efficient MoM estimator with low sample complexity remains an open and important direction for future research.
@@ 5.2 Simulation studies
Appendix A contains the full comparison table with existing theoretical works of EM. Lemma 1 is proved in Appendix B. The proofs of population-level convergence of the EM algorithm are stated in Appendix C. Appendix D contains the proofs of the sample-level results in Section 4.1 when the separation is small to moderate while Appendix E contains the proofs when the separation is moderate to large. Appendix F contains the proof of the minimax lower bounds. Auxiliary lemmas are collected in Appendix G.
@@ Appendix A A full comparison table with existing theoretical works of the EM algorithm under the $L$-GMM
Table 3: Comparison with existing theoretical results for the EM algorithm under the $L$-GMM. Symmetric 2-GMM means $\pi_{1}=\pi_{2}=1/2$ and $\mu_{1}=-\mu_{2}=\mu$. In [cai2019chime], $\beta=\Sigma^{-1}(\mu_{2}-\mu_{1})$.
@@
[EM2017] 22 symmetric, $\Sigma=\mathbf{I}_{d}$ >C>C $\|\mu^{(0)}-\mu\|_{2}\leq\|\mu\|_{2}/4$ $d(\widehat{M},M)\lessapprox\|\mu\|_{2}^{3}\sqrt{d/n})$
[Xu_Hsu_Maleki] 22 symmetric, $\Sigma$ known >0>0 $\mu^{(0)\top}\mu\neq 0$ $d(\widehat{M},M)=o(1)$
[daskalakis2017ten] 22 symmetric, $\Sigma$ known $\widetilde{\Omega}((d/n)^{1/3})$ $\mu^{(0)\top}\mu\neq 0$ $d(\widehat{M},M)\lessapprox(\sqrt{d/n}\wedge\|\mu\|_{2}^{-2}\sqrt{d/n})$
[wu2021randomly] 22 symmetric, $\Sigma=\mathbf{I}_{d}$ $(0,C]$ $\mu^{(0)\top}\mu\neq 0$ $d(\widehat{M},M)\lessapprox\left((d/n)^{1/4}\wedge\|\mu\|_{2}^{-1}\sqrt{d/n}\right)$
[weinberger2022algorithm] 22
@@
$(0,C]$ $\mu^{(0)\top}\mu\geq 0$ $d(\widehat{M},M)\lessapprox\left((d/n)^{1/4}\wedge\|\mu\|_{2}^{-1}\sqrt{d/n}\wedge(\pi_{2}-\pi_{1})^{-1}\sqrt{d/n}\right)$
[cai2019chime] 22
@@
$d(\widehat{M},M)+\|\widehat{\beta}-\beta\|_{2}\lesssim\sqrt{d/n}$
[dasgupta2007probabilistic] $\geq 2$ $\Sigma=\sigma^{2}\mathbf{I}_{d}$ $\Omega\Bigl(\sqrt{d\log\pi_{\min}^{-1}}+\log\pi_{\min}^{-1}\Bigr)$ A particular initialization $d(\widehat{\pi},\pi)+d(\widehat{M},M)\lesssim\left(\sqrt{d/(n\pi_{\min})}+{\exp(-c\sqrt{d})/\pi_{\min}}\right)$
[kwon2020algorithm] $\geq 2$ $\Sigma=\sigma^{2}\mathbf{I}_{d}$ $\log(1/\pi_{\min})$ $d(\pi^{(0)},\pi)<1,~d(M^{(0)},M)\leq c\sqrt{\Delta_{\min}}$ A sample-splitting version of the EM
[yan2017convergence] $\geq 2$ $\pi$ known, $\Sigma=\mathbf{I}_{d}$ $\Omega\Bigl((d\wedge L)\log{(\Delta_{\max}/\pi_{\min})}\Bigr)$ $d(M^{(0)},M)\leq c\sqrt{\Delta_{\min}}$ $d(\widehat{M},M)\lessapprox{(\pi_{\max}/\pi_{\min})}\sqrt{\Delta_{\max}}(\sqrt{d}+L^{3}\Delta_{\max})\sqrt{d/n}$
[Zhao2020] $\geq 2$ $\pi$ known, $\Sigma=\mathbf{I}_{d}$ $\Omega\Bigl((d\wedge L)\log{(\Delta_{\max}/\pi_{\min})}\Bigr)$ $d(M^{(0)},M)\leq c\sqrt{\Delta_{\min}}$ $d(\widehat{M},M)\lessapprox({\sqrt{\Delta_{\max}}/\pi_{\min}})\sqrt{dL/n}$
[segol2021improved] $\geq 2$ $\pi$ known, $\Sigma=\mathbf{I}_{d}$ $\log(1/\pi_{\min})$ $d(M^{(0)},M)\leq c\sqrt{\Delta_{\min}}$ $d(\widehat{M},M)\lessapprox(1/\pi_{\min})\sqrt{dL/n}$
This work $\geq 2$ Full generality $\Omega\Bigl(\log{(\Delta_{\max}/\pi_{\min})}\Bigr)$ $d(\pi^{(0)},\pi)<1,~d(M^{(0)},M)\leq c\sqrt{\Delta_{\min}}$ $d(\widehat{\pi},\pi)+d(\widehat{M},M)\lessapprox\sqrt{d/(n\pi_{\min})}$
$d(\Sigma^{(0)},\Sigma)\leq c$, or via preliminary labels
$d(\widehat{\Sigma},\Sigma)\lessapprox\sqrt{d/n}$
@@ C.3 Proof of Theorem 2: contraction of the population level EM iterates
$\begin{split}\mathcal{R}_{I}^{(0)}&:=\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\max_{\ell\in[L]}\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)\right],\\ \mathcal{R}_{I}^{(1)}&:=\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\max_{\ell\in[L]}~\Bigl\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)NN^{\top}\right]\Bigr\|_{\rm op},\\ \mathcal{R}_{I}^{(2)}&:=\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\max_{\ell\in[L]}\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\left\|\mathbb{E}\left[I_{a\ell}(X,\omega)NN^{\top}\right]\right\|_{\rm op},\end{split}$ (C.22)
@@
(C.36) $\leq\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}[I_{a\ell}(X,\omega)]\max_{b\in[L]}\left(\sqrt{\Delta_{b\ell}}+\sqrt{\Delta_{a\ell}}\right)\left\|\Sigma^{*1/2}(\bar{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{2}$
@@
For any $\omega=(\pi,M,J)$ and $\ell\in[L]$, note that
$\bm{\mu}_{\ell}(\omega)-\mu_{\ell}^{*}=\frac{\mathbb{E}[\gamma_{\ell}(X;\omega)(X-\mu_{\ell}^{*})]}{\mathbb{E}[\gamma_{\ell}(X;\omega)]}={1\over\bm{\pi}_{\ell}(\omega)}\mathbb{E}\left[(\gamma_{\ell}(X;\omega)-\gamma_{\ell}(X;\omega^{*}))(X-\mu_{\ell}^{*})\right].$
Fix any $\bar{\omega}\in\mathcal{B}_{d}(\omega^{*})$ and any $\ell\in[L]$. It then follows that
@@
$\begin{split}\nabla^{\pi}\bm{\mu}_{\ell}(\bar{\omega})&:=\sum_{k=1}^{L-1}\int_{0}^{1}\frac{1}{\bm{\pi}_{\ell}(\omega)}\mathbb{E}\left[\left(X-\mu_{\ell}^{*}\right)\left\langle{\partial\gamma_{\ell}(X;\omega)\over\partial\pi_{k}}\Big|_{\omega=(\pi_{u},\bar{M},\bar{J})},~\left(\bar{\pi}_{k}-\pi_{k}^{*}\right)\right\rangle\right]{\rm d}u\\ \nabla^{M}\bm{\mu}_{\ell}(\bar{\omega})&:=\sum_{k=1}^{L}\int_{0}^{1}\frac{1}{\bm{\pi}_{\ell}(\omega)}\mathbb{E}\left[\left(X-\mu_{\ell}^{*}\right)\left\langle{\partial\gamma_{\ell}(X;\omega)\over\partial\mu_{k}}\Big|_{\omega=(\pi^{*},M_{u},\bar{J})},~\bar{\mu}_{k}-\mu_{k}^{*}\right\rangle\right]{\rm d}u\\ \nabla^{J}\bm{\mu}_{\ell}(\bar{\omega})&:=\int_{0}^{1}\frac{1}{\bm{\pi}_{\ell}(\omega)}\mathbb{E}\left[\left(X-\mu_{\ell}^{*}\right)\left\langle{\partial\gamma_{\ell}(X;\omega)\over\partial J}\Big|_{\omega=(\pi^{*},M^{*},J_{u})},~\bar{J}-J^{*}\right\rangle\right]{\rm d}u.\end{split}$ (C.42)
The main task below is to bound each terms in terms of $\mathcal{R}_{I}^{(j)}$ with $j\in\{0,1,2\}$. For future reference, we notice that
@@
$\left\|\nabla^{\pi}\bm{\mu}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}\leq\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\frac{1}{\bm{\pi}_{\ell}(\omega)}\left\|\sum_{k=1}^{L-1}(\bar{\pi}_{k}-\pi_{k}^{*})~\mathbb{E}\left[\left(X-\mu_{\ell}^{*}\right){\partial\gamma_{\ell}(X;\omega)\over\partial\pi_{k}}\right]\right\|_{\Sigma^{*}}.$